# Mind the Gap. Doubling Constant Parametrization of Weighted Problems: TSP, Max-Cut, and More

### v1

@@ 4.4 Steiner Meets Freiman
For reference, we also include the Minimum Steiner Tree, even though the general setting can be directly solved in $O((2+\epsilon)^{k}\,p(n))$ [faster_steiner1], where $p(n)$ is a polynomial function of $n$, the degree of which grows rapidly as $\epsilon$ approaches zero, Its degree has been later refined to $12\sqrt{\epsilon^{-1}\ln\epsilon^{-1}}$, having as a result running times such as $O(2.1^{k}\,n^{57.6})$ or $O(2.5^{k}\,n^{14.2})$ [fuchs2007dynamic]. Notably, Dreyfus and Wagner [dreyfus1971steiner] were the first to design a dynamic programming recursion running in time $O(3^{k}n+2^{k}n^{2}+nm)$, where $n=|V|$, $m=|E|$, and $k=|T|$.
@@
Condition 1 holds since, as the edge weights are all non-negative, the optimal subgraph $H$ is a tree with its leaves in $K$, and a tree’s weight is the sum of the weights of its constituent edges. Next, we can take $\mathcal{A}$ as the fast subset convolution based algorithm due to Björklund, Husfeldt, Kaski, and Koivisto [fsc, Sec. 4.1.2], which runs in $O^{*}(2^{k}n^{2}W)$-time, or Lokshtanov–Nederlof’s polynomial-space algorithm [saving_space] in the same running time. ∎
@@ 5 Outlook: Beyond NP-hard Problems
The answer seems to be a positive one, in particular for those polynomial problems for which the solution is an additive combination of the input weights. A simple application is that of the $(\min,+)$-convolution: Given sequences $(a[i])_{i=0}^{n-1}$ and $(b[i])_{i=0}^{n-1}$, the goal is to compute $(c[i])_{i=0}^{n-1}$, where $c[k]=\min_{i=0,\ldots,k}\{a[i]\,+\,b[k-i]\}$. The naive algorithm runs in $O(n^{2})$-time, and no $O(n^{2-\varepsilon})$-time algorithm is known for $\varepsilon>0$ [min-plus-1, min-plus-2]. However, the convolution in the $(+,\times)$-ring can be solved in $O(n\log n)$-time via FFT. This resembles the setting we considered throughout the paper. Indeed, Węgrzycki outlines in his PhD thesis a bounded-input algorithm running in time $\widetilde{O}(nW)$, where $W$ is the largest input value [karol_phd, Lemma 5.7.2]. Applying the same key idea as for the NP-hard problems we considered, we can obtain an $\widetilde{O}_{\mathcal{C}}(n)$-time algorithm for the min-plus self-convolution problem, i.e., $a=b$, and the input sequence, when regarded as a set, has small doubling. The same setting holds true for the min-sum subset convolution [fsc].
It is thus interesting to ask how the framework behaves in more weighted applications, such as (specialized) $(\min,+)$ matrix products [special-min-plus-prod] or even APSP, where—similar to the NP-hard case studied here—the objective value is an additive combination of input weights.
@@ Appendix B Edge-Weighted $k$-Clique with Small Weights
We show that Nešetřil-Poljak’s algorithm [uw-k-clique] can be extended to support small weights. We consider the $3k$-case here; both the $3k+1$- and $3k+2$-cases are handled analogously.

### v2

@@ 4.4 Steiner Meets Freiman
For reference, we also include the Minimum Steiner Tree, even though the general setting can be directly solved in $O((2+\epsilon)^{k}\,p(n))$ [MRR06], where $p(n)$ is a polynomial function of $n$, the degree of which grows rapidly as $\epsilon$ approaches zero, Its degree has been later refined to $12\sqrt{\epsilon^{-1}\ln\epsilon^{-1}}$, having as a result running times such as $O(2.1^{k}\,n^{57.6})$ or $O(2.5^{k}\,n^{14.2})$ [FKM+07]. Notably, Dreyfus and Wagner [DW71] were the first to design a dynamic programming recursion running in time $O(3^{k}n+2^{k}n^{2}+nm)$, where $n=|V|$, $m=|E|$, and $k=|T|$.
@@
Condition 1 holds since, as the edge weights are all non-negative, the optimal subgraph $H$ is a tree with its leaves in $K$, and a tree’s weight is the sum of the weights of its constituent edges. Next, we can take $\mathcal{A}$ as the fast subset convolution based algorithm due to Björklund, Husfeldt, Kaski, and Koivisto [BHKK07, Sec. 4.1.2], which runs in $O^{*}(2^{k}n^{2}W)$-time, or Lokshtanov–Nederlof’s polynomial-space algorithm [LN10] in the same running time. ∎
@@ 5 Outlook: Beyond NP-hard Problems
The answer seems to be a positive one, in particular for those polynomial problems for which the solution is an additive combination of the input weights. A simple application is that of the $(\min,+)$-convolution: Given sequences $(a[i])_{i=0}^{n-1}$ and $(b[i])_{i=0}^{n-1}$, the goal is to compute $(c[i])_{i=0}^{n-1}$, where $c[k]=\min_{i=0,\ldots,k}\{a[i]\,+\,b[k-i]\}$. The naive algorithm runs in $O(n^{2})$-time, and no $O(n^{2-\varepsilon})$-time algorithm is known for $\varepsilon>0$ [CMWW17, KPS17]. However, the convolution in the $(+,\times)$-ring can be solved in $O(n\log n)$-time via FFT. This resembles the setting we considered throughout the paper. Indeed, Węgrzycki outlines in his PhD thesis a bounded-input algorithm running in time $\widetilde{O}(nW)$, where $W$ is the largest input value [Węg19, Lemma 5.7.2]. Applying the same key idea as for the NP-hard problems we considered, we can obtain an $\widetilde{O}_{\mathcal{C}}(n)$-time algorithm for the min-plus self-convolution problem, i.e., $a=b$, and the input sequence, when regarded as a set, has small doubling. The same setting holds true for the min-sum subset convolution [BHKK07].
It is thus interesting to ask how the framework behaves in more weighted applications, such as (specialized) $(\min,+)$ matrix products [VX20] or even APSP, where—similar to the NP-hard case studied here—the objective value is an additive combination of input weights.
@@ Appendix B Edge-Weighted $k$-Clique with Small Weights
We show that Nešetřil-Poljak’s algorithm [NP85] can be extended to support small weights. We consider the $3k$-case here; both the $3k+1$- and $3k+2$-cases are handled analogously.